The smaller of two numbers is eight less than three times the larger number. If twice the smaller number is added to 15, the result is 11. Find each of the numbers. Please help me!!
Answer by lwsshak3(11628)

The smaller of two numbers is eight less than three times the larger number. If twice the smaller number is added to 15, the result is 11. Find each of the numbers.
**
let x = smaller of two numbers
let y = larger of two numbers
..
x=3y-8
2x+15=11
2(3y-8)+15=11
6y-16+15=11
6y=12
y=2
x=3y-8=-2
Ans:
smaller of two numbers=-2
larger of two numbers=2